how many milimeters of a 15% acid solution and how many milimeters of 3% acid solution must be mixed to get 180 milimeters of a 10% acid solution
Answer by mananth(16946)

---------percent ---------------- quantity
Acid I 15 ---------------- x ml
Acid II 3 ---------------- 120 - x ml
Mixture 10.00% ---------------- 120

15 x + 3 ( 120 - x ) = 10 * 120

15x+ 360 - 3 x = 1200
15 x - 3 x = 1200 - 360
12 x = 840
/ 12
x=70 ml 15.00% Acid I
50 ml 3.00% Acid II